write the first five terms of the sequence in which the nth term is 5n-4
Answer by Edwin McCravy(20054)

write the first five terms of the sequence in which the nth term is 5n-4
1. Substitute 1 for n in 5n-4. You get 5(1)-4 = 5-4 = 1, so the 1st term is 1.

2. Substitute 2 for n in 5n-4. You get 5(2)-4 = 10-4 = 6, so the 2nd term is 6.

3. Substitute 3 for n in 5n-4. You get 5(3)-4 = 15-4 = 11, so the 3rd term is 11.

4. Substitute 4 for n in 5n-4. You get 5(4)-4 = 20-4 = 16, so the 4th term is 16.

5. Substitute 5 for n in 5n-4. You get 5(5)-4 = 25-4 = 21, so the 5th term is 21.

Now list those 5 terms in a row:

1, 6, 11, 16, 21.

Don't you wish all your problems were this easy?

Edwin